OpenStudy (anonymous):

use radical notation to write the expression: 5m^1/3

OpenStudy (anonymous):

\[\sqrt[3]{5m}\]

jimthompson5910 (jim_thompson5910):

\[\large 5m^{\frac{1}{3}}=5\sqrt[3]{m}\]

OpenStudy (anonymous):

two different answers... ??

OpenStudy (anonymous):

Yea i guess so i like the one i put and i'm sure jim likes his so your going to have to pick i guess

jimthompson5910 (jim_thompson5910):

If the original expression is \[\large 5m^{\frac{1}{3}}\], then the answer I posted is correct

jimthompson5910 (jim_thompson5910):

Or, if the original expression is \[\large (5m)^{\frac{1}{3}}\], then the other answer is correct. See the difference?

jimthompson5910 (jim_thompson5910):

It all depends on whether or not there are parenthesis surrounding the "5m"